# Detecting Multimedia Generated by Large AI Models: A Survey

### v1

@@ 3.1.1 Pure Detection
Watermarking. Detecting watermarks in the text would be the most straightforward way for humans to understand the detection technique. Text watermarking injects algorithmically detectable patterns into the generated text while ideally preserving the quality and diversity of language model outputs. Most detection methods based on watermark require the watermark key for detection [212, 213, 214, 215, 216, 217], which are susceptible to security breaches and counterfeiting. [218] proposes the first method, which does not require the key during detection to alleviate this issue. They use the shared token embedding between the watermark generation and detection networks to improve the efficiency of training the detector.
Artifacts. While artifacts in text may not be as visually obvious as in images, this detection approach is simple and comprehensible to humans. Yet, intriguing artifacts or characteristics of LLM-generated texts can be elusive. [211] reveals those characteristics by perturbing these generated texts and offers the following empirical findings:
  • Artifacts exist mainly in the form of token co-occurrence and at the head of the vocabulary;
  • Content words contain more artifacts than stopwords;
  • High-level semantic/syntactic features contain much fewer artifacts than shallow features;
  • Some artifacts are present in higher-order N-grams, appearing as elements with unclear or vague meanings.
Stylometry/Coherence. Other works directly identify the stylistic changes and coherence inconsistency between LLM-generated texts and human-written ones. [209] employs stylometric analysis to distinguish between human and LLM-generated tweets by quantifying stylistic differences. [210] examines the linguistic structure of texts and uses the difference in coherence traced by entity consistency exhibited in LLM-generated text and human-written ones.
$\clubsuit$ Hard-explainable Methods. These methods involve detection techniques and analytical processes that may not be as easily comprehensible to most people but are quite accessible to researchers in the field.
Perplexity. This is a statistical metric used in language models. Several works employ perplexity to distinguish LLM-generated text from human-written text. For example, [207] discerns LLM-generated text, specifically homework assignments, by computing perplexity scores for student-authored and ChatGPT-generated responses. These scores then assist in establishing a threshold for discerning the origin of a submitted assignment. Moreover, the widely recognized tool, GPTZero [208], examines the text’s perplexity and burstiness metrics to estimate the likelihood of a review text being generated by LLMs for detection.
Log Probabilities Curvature. This refers to the logarithmic transformation of probabilities that a model assigns to sequences of words or tokens, and be used in certain research to detect LLM-generated text. Some works statistically utilize log probabilities to detect LLM-generated text, which require a certain level of knowledge among the general public regarding the generative mechanisms of LLMs to understand their detection methods. For example, the pioneer work DetectGPT [206] observes that LLM-generated text exhibits a more negative log probability curvature and leverages the curvature-based criterion based on random perturbations of the text, yielding promising results.
Efficient Perturbations. Recognizing the use of pure random perturbations in DetectGPT requires intensive computational cost, a series of works adopt various techniques to improve computational efficiency and enhance detection accuracy simultaneously. Specifically, [203] achieves similar performance with up to two times fewer queries than DetectGPT with a Bayesian surrogate model by selecting typical samples based on Bayesian uncertainty and interpolating scores from typical samples to other ones, making the perturbation process more focused and less resource-intensive. Bao et al. [204] increase the detection speed by 340 times by substituting DetectGPT’s perturbation step with a more efficient sampling step via conditional probability curvature. DetectLLM [205], another recent contribution, employs normalized perturbed log-rank for text detection generated by LLMs, asserting a lower susceptibility to the perturbation model and the number of perturbations compared to DetectGPT.
Positive-Unlabeled. Tian et al. [202] observe that LLM-generated short texts are highly similar to human, it is not suitable to assign these simple generated texts with either fake or real labels; rather, they are in an “Unlabeled” state. To this end, they model the task of LLM-generated text detection as a partial Positive-Unlabeled problem and formulate the Multiscale Positive-Unlabeled training framework to address the challenging task of short text detection without sacrificing long texts.
@@ 3.1.2 Beyond Detection
$\clubsuit$ Attribution. Determining which specific model may generate the test content, see Fig. 5 a).
Deep-learning Based. Most attribution models are deep-learning based [198, 199, 200, 201]. [200] addresses attribution by noting that synthetic texts carry subtle distinguishing marks inherited from their source models and these marks can be leveraged for attribution. TopRoBERTa [201] improves existing attribution solutions by capturing more linguistic patterns in LLM-generated texts by including a Topological Data Analysis (TDA) layer in the RoBERTa model. In this approach, RoBERTa RoBERTa captures contextual semantic and syntactic features, while a TDA layer analyzes the shape and structure of linguistic data.
Stylometric/Statistical. Besides the above deep learning-based classifier, some works solve the attribution task in a stylometric or statistical way. Stylometry is the statistical analysis of writing styles in texts. Uchendu et al. [29] propose a model named the Linguistic model. They train a Random Forest classifier with their proposed Authorship Attribution (AA) dataset and extract an ensemble of stylometric features, such as entropy and readability score. GPT-who [197] is a psycholinguistically-aware, statistical-based detector. It utilizes Uniform Information Density (UID) based features to model the unique statistical signature of each LLM and human author for accurate authorship attribution.
Perplexity. Inspired by the observation that perplexity serves as a reliable signal for distinguishing the source of generated text, [196] calculates the proxy perplexity to identify the sources from which the text is generated, such as Human, LLaMA, OPT [219], or others.
Style Representation. In pursuit of an approach that does not rely on samples from language models of concern at training time. Soto et al. [195] propose a few-shot strategy to detect LLM-generated text by leveraging representations of writing style estimated from human-written text.
Origin Tracing. This differs from the above model-wise attribution. It refers to tracing back a larger original model on which a smaller generative model is based. For example, trace text generated by Alpaca [220] back to ChatGPT and LLaMA. [194] introduces the first origin tracing tool, Sniffer, which utilizes the discrepancies between LLMs as features and then trains a simple linear classifier to help trace the origins.
$\clubsuit$ Generalization. Developing detectors with generalizability that can detect texts generated by generators never seen before, as shown in Fig. 5 b).
Structured Search. This method involves passing documents through weaker language models and then conducting a systematic search over possible combinations of their features. Verma et al. present Ghostbuster [23], a method for detection with generalizability based on structured search and linear classification. Ghostbuster runs the structured search to get the possible text features and then trains a classifier on the selected features to predict whether documents are AI-generated.
Contrastive Learning. [193] develops a contrastive domain adaption framework that blends standard domain adaption techniques with the representation power of contrastive learning to learn domain invariant representations for better generalizability.
$\clubsuit$ Interpretability. Exploring interpretable detectors that can provide explanations for detection results, see Fig. 5 c).
N-gram Overlaps. DNA-GPT [192] identifies GPT-generated text by exploiting the distinct text continuation patterns between human and AI-generated content. It provides evidence based on nontrivial N-gram overlaps to support explainable detection results.
P-values. [191] proposes a statistical test for detecting watermarks in text with interpretable P-values. P-values provide a statistically interpretable way to quantify the certainty that the detected pattern (i.e., watermark) in the text is not due to random chance.
Shapley Additive Explanations. Some researchers work toward gaining insight into the reasoning behind the model. For example, [190] fine-tunes a transformer-based model and uses it to make predictions, which are then explained using Shapley Additive Explanations (SHAP) [221], an eXplainable Artificial Intelligence (XAI) framework to extract explanations of the model’s decisions aiming at uncovering the model’s reasoning. Similarly, [38] also employs Shapley Values to compare the interpretations derived from word-level and sentence-level results.
Polish Ratio. Though the above studies provide interpretability, they do not address text that has been refined on a more granular level, such as ChatGPT-polished texts. To bridge this gap, [39] introduces a novel dataset termed HPPT (ChatGPT-polished academic abstracts) and proposes the “Polish Ratio" method, a measure of ChatGPT’s involvement in text based on editing distance. It calculates a similarity score between the original human-written texts and their polished versions, providing a more reasonable explanation for the detection outcome.
@@
$\clubsuit$ Robustness. Developing detectors that can handle different attacks, see Fig. 5 d). In light of the vulnerability of detectors to different attacks and robustness issues, a significant body of research has been dedicated to utilizing adversarial learning as a mitigation strategy.
Adversarial Data Augmentation. [39, 188, 189] conduct the adversarial data augmentation process on LLM-generated text, their findings indicate that models trained on meticulously augmented data exhibited commendable robustness against potential attacks. The key technique in these methods involves employing adversarial attacks by creating misleading inputs, thereby enhancing the model’s competency to handle a wider range of situations where deception might be a factor.
Adversarial Learning. Besides these pre-processing methods, some works cast the spotlight on adversarial learning-based frameworks. Hu et al. [187] introduce RADAR, a framework for training robust detectors by simultaneously interacting with a paraphrasing model, which generates content aimed at evading detection. While effective against paraphrase attacks, its defense against other attacks is unexplored. Diverging from RADAR, Koike et al. [32] propose OUTFOX, which enhances robustness by interlinking detector and attacker outputs. The attacker learns from the detector’s predictions to create hard-to-detect essays, while the detector improves by learning from these adversarial essays. OUTFOX is particularly effective against paraphrase-based attacks.
Stylistic/Consistency. [185] studies and quantifies stylistic cues from the latent journalism process in real-world news organizations towards discriminating AI-generated news. Their J-Guard framework steers existing supervised AI text detectors while boosting adversarial robustness. The research by Tulchinskii et al. [186] explores consistent properties in human-written texts across different domains and skill levels. They estimate the geometry of text sample as an individual object and discover that real texts have a higher intrinsic dimension than artificial ones. This insight is used to estimate intrinsic dimensionality for detecting LLM-generated texts.
$\clubsuit$ Empirical Study. The empirical studies are crucial for advancing our understanding and capabilities in detecting LLM-generated texts, as shown in Fig. 5 e).
Generalization/Robustness. Tu et al. [182] observe a decline in RoBERTa-based detectors’ effectiveness over time through a month-long study of ChatGPT responses to long-form questions. Complementing this, Pu et al. [183] find that detectors trained on one generator can zero-shot generalize to another, especially when trained on a medium-size LLM for detecting content from a larger version.
A more comprehensive investigation on detectors’ generalization is conducted by Xu et al. [48]. They find that the trained models tend to overfit to some “irrelevant features” that are not principal for ChatGPT detection. This overfitting issue can be originated from the “incautious and insufficient” data collection process. Besides this, they provide an optimistic insight: the trained models are also capable of extracting “transferable features”, which are shared features that can help detect the ChatGPT generated texts from various topics and language tasks.
Additionally, [184] specifically investigate generalization of in-domain fine-tune detectors. They use the multi-domain, multilingual AuTexTification corpus to fine-tune various supervised detectors and discover that in-domain fine-tune detectors struggle against data from different models. Moreover, detection in the wild [44] conducts a more comprehensive experiment by building a wild testbed, involving 10 datasets covering diverse writing tasks and sources, and using 27 LLMs for creating texts. A key finding is the challenge of Out-Of-Distribution data in real-world scenarios, where the performance of detectors significantly declines, often barely better than random classification.
Human Evaluation. [50] conducts thorough human evaluations and linguistic analyses to compare the content generated by ChatGPT with that produced by humans. Key findings include: ChatGPT’s responses are more objective and formal, less likely to contain biased or harmful information, and generally longer and more detailed.
Attribution. An extensive investigation on model attribution, encompassing source model identification, model family classification, and model size classification, is conducted by Antoun et al. [181]. Their source model identification task involves classifying the text into one of the 50 LLMs, spanning various families and sizes. The results reveal several key findings: a clear inverse relationship between classifier effectiveness and model size, with larger LLMs being more challenging to detect, especially when the classifier is trained on data from smaller models.
Paraphrase Detection. Beyond detecting texts purely generated by LLMs and written by humans. Another line of empirical studies [27, 180] conduct extensive experiments on paraphrase detection concerning the severe threat to academic integrity. Specifically, [27] evaluates the performance of various detectors and perform a human study with 105 participants regarding their detection performance and the quality of generated examples. Their results suggest that humans struggle to identify large models-paraphrased text (53% mean accuracy). Human experts rate the quality of paraphrases generated by GPT-3 to be as high as that of original texts (clarity 4.0/5, fluency 4.2/5, coherence 3.8/5).  [180] evaluates various detection methods on several common paraphrase datasets, finding that human-authored paraphrases exceed machine-generated ones in terms of difficulty, diversity, and similarity.

### v4

@@ 3.1.1 Pure Detection
Watermarking. One intuitive direction is text watermarking, where detectable patterns are algorithmically embedded in generated content. While most watermarking methods require a secret key for detection [330, 331, 332, 333, 334, 335], Liu et al. [336] address this limitation by introducing a keyless approach using shared token embeddings between the generator and detector.
Non-watermarking. Another line of work explores textual artifacts—subtle irregularities in token usage or structure left by language models. Pu et al. [325] reveal that artifacts tend to appear in token co-occurrence patterns, especially within the head of the vocabulary and in shallow features like high-frequency content words or higher-order N-grams.
Several studies also rely on stylistic or coherence cues to differentiate human from LLM-generated text. Kumarage et al.[326] apply stylometric analysis to quantify differences in writing style, while Liu et al.[327] assess entity-level coherence to expose inconsistency patterns more common in machine-generated text.
Finally, Zhu et al. [328] propose a zero-shot, black-box method that queries ChatGPT to revise input text. The method assumes ChatGPT makes fewer edits to AI-generated text and detects generation by measuring similarity between the original and revised versions—without needing model logits or additional training. Wu et al. [329] propose GECScore, a zero-shot, black-box method that detects LLM-generated text by measuring similarity between the input and its grammar-corrected version. Based on the assumption that LLM text contains fewer grammatical errors, this approach offers a simple detection signal.
$\clubsuit$ Hard-explainable Methods. These methods involve detection techniques and analytical processes that may not be as easily comprehensible to most people but are quite accessible to researchers in the field. Several hard-explainable methods leverage statistical signals or weak supervision to detect LLM-generated text. Perplexity, a standard metric in language modeling, is widely used to distinguish between human- and machine-generated content. For instance, [320] computes perplexity scores to separate student-written from ChatGPT-generated homework submissions, while GPTZero [321] combines perplexity and variance in sentence-level likelihood to assess the likelihood of text being AI-generated. Hans et al.[322] further propose Binoculars, a zero-shot detector that uses the ratio of perplexity to cross-perplexity between two LLMs, offering a strong, training-free signal for detection across domains. Similarly, log probability curvature has been explored in DetectGPT[323], which observes that LLM-generated text exhibits more negative curvature in the log probability landscape. By perturbing the input and analyzing changes in curvature, DetectGPT achieves promising zero-shot results. Beyond statistical modeling, Positive-Unlabeled (PU) training has also been applied to tackle label uncertainty. Tian et al. [324] argue that short generated texts often resemble human-written content closely and thus are better treated as unlabeled. They formulate LLM-generated text detection as a PU learning problem and propose a multiscale PU framework that maintains detection accuracy across text lengths.
@@ 3.1.2 Beyond Detection
$\clubsuit$ Efficiency. Recognizing the use of pure random perturbations in DetectGPT [323] requires intensive computational cost, a series of works adopt various techniques to improve computational efficiency and enhance detection accuracy simultaneously, as shown in Fig. 5 a). Specifically, [315] achieves similar performance with up to two times fewer queries than DetectGPT with a Bayesian surrogate model by selecting typical samples based on Bayesian uncertainty and interpolating scores from typical samples to other ones, making the perturbation process more focused and less resource intensive. Bao et al. [316] increase the detection speed by 340 times by substituting DetectGPT’s perturbation step with a more efficient sampling step via conditional probability curvature. DetectLLM [317], another recent contribution, employs normalized perturbed log-rank for text detection generated by LLMs, asserting a lower susceptibility to the perturbation model and the number of perturbations compared to DetectGPT. SeqXGPT [318] enhances the efficiency of log probability-based detection by operating at the sentence level without requiring costly perturbations. It leverages token-level log probabilities from multiple white-box LLMs and processes them using lightweight convolutional and self-attention networks to capture nuanced statistical patterns that effectively separate human-written from machine-generated sentences. Glimpse [319] significantly improves runtime and cost-efficiency in zero-shot detection by extending log-probability curvature methods to proprietary LLMs. It estimates full token distributions from limited top-K API outputs, enabling accurate detection with substantially fewer API calls and reduced computational overhead, even for powerful models like GPT-4.
$\clubsuit$ Attribution. Determining which specific model may generate the test content, see Fig. 5 b). Many recent approaches employ deep learning models to capture semantic and structural signals embedded in the text. For instance, Munir et al.[313] highlight that synthetic texts carry subtle, consistent traces from their originating models, which can be leveraged for attribution. TopRoBERTa[314] enhances attribution performance by integrating topological data analysis (TDA) with RoBERTa, enriching contextual semantic and syntactic representations with structural insights. Other methods rely on explicitly defined statistical and linguistic features. Uchendu et al.[168] use a Random Forest classifier trained on stylometric indicators such as entropy and readability. GPT-who[307] introduces Uniform Information Density (UID)-based psycholinguistic features to distinguish human from model-generated text, while Wu et al.[308] utilize proxy perplexity scores for attributing text to models like LLaMA and OPT[337]. Soto et al.[309] take a few-shot approach, estimating style representations solely from human-authored samples to avoid dependence on generated data. Moving beyond attribution to a single model, Sniffer[310] infers generative lineage by analyzing statistical discrepancies, enabling traceability across models, for example, linking Alpaca [338] back to ChatGPT and LLaMA.
$\clubsuit$ Generalization. Developing detectors with generalizability that can detect texts generated by generators never seen before, as shown in Fig. 5 c). [37], which first uses weaker language models to extract structured features from texts through systematic search, and then trains a linear classifier on these selected features to detect AI-generated content. Bhattacharjee et al. [303] propose a contrastive domain adaptation framework, combining traditional domain adaptation techniques with contrastive learning to learn domain-invariant representations, thereby improving the generalization ability of detectors. Text Fluoroscopy [304] enhances generalization by leveraging intrinsic features from intermediate layers of language models, selecting the most distinctive layer to enable improved cross-domain detection. DeTeCtive [305] further advances this line of work by introducing a multilevel contrastive learning approach that treats detection as a style discrimination problem. It encodes fine-grained stylistic features using auxiliary contrastive objectives, achieving strong generalization to out-of-distribution generators. Notably, it supports training-free incremental adaptation, enabling detectors to incorporate new texts during inference without retraining. Tulchinskii et al. [306] propose a generator-agnostic method based on intrinsic dimensionality (ID) estimation of sentence embeddings. Without any training, their method generalizes across unseen domains, generators, and languages.
$\clubsuit$ Interpretability. Exploring interpretable detectors that can provide explanations for detection results, see Fig. 5 d). DNA-GPT[300] introduces a training-free approach that identifies GPT-generated text by analyzing distinct text continuation patterns between human and AI-generated content, offering evidence based on nontrivial N-gram overlaps to support explainable detection results. Kirchenbauer et al.[301] propose a statistical test for detecting watermarks in text, utilizing interpretable p-values to quantify the certainty that the detected pattern (e.g., watermark) is not due to random chance. To gain insight into model reasoning, Mitrovic et al.[302] fine-tune a transformer-based model and employ Shapley Additive Explanations (SHAP) [339] to extract explanations of the model’s decisions, aiming to uncover the model’s reasoning. Similarly, Liu et al.[60] utilize Shapley Values to compare interpretations derived from word-level and sentence-level results. Addressing the challenge of detecting texts refined at a granular level, such as ChatGPT-polished texts, Yang et al.[46] introduce a novel dataset termed HPPT (ChatGPT-polished academic abstracts) and propose the "Polish Ratio" method, which measures ChatGPT’s involvement in text based on editing distance, providing a more reasonable explanation for the detection outcome
@@
$\clubsuit$ Robustness. Developing detectors that can handle different attacks, see Fig. 5 e). In light of the vulnerability of detectors to different attacks and robustness issues, a significant body of research has been dedicated to utilizing adversarial learning as a mitigation strategy.
Adversarial Attack Robustness. A growing body of work has focused on improving the robustness of AI-generated text detectors against adversarial perturbations, including paraphrasing, synonym substitution, token-level manipulations, and stylistic shifts. RADAR [298] introduces a detector trained in tandem with a paraphrasing adversary via adversarial learning, significantly improving detection robustness under paraphrased attacks. Similarly, OUTFOX [41] advances this idea by co-evolving a detector and an attacker through in-context learning, enabling the detector to adapt to increasingly sophisticated adversarial examples. Red teaming approaches [297] further stress-test detection systems by using LLMs to generate synonym-based or instruction-prompted adversarial variants, revealing that even strong detectors can be bypassed when facing semantically equivalent but stylistically modified text. J-Guard [299], in contrast, enhances adversarial robustness by encoding journalism-specific stylistic features, such as AP-style conventions, which are less susceptible to word-level transformations. Together, these works demonstrate that models explicitly trained to handle adversarial perturbations, particularly through dynamic attacker-detector interactions or domain-guided cues, are better equipped to maintain detection reliability in real-world settings where malicious evasion is increasingly likely.
LAIM-Polished Robustness. Yang et al. [46] address the challenge of detecting partially AI-influenced text, such as human-written content polished by ChatGPT. They introduce the HPPT dataset and propose the Polish Ratio to quantify ChatGPT’s editing impact, showing improved robustness in identifying collaborative writing scenarios.
$\clubsuit$ Empirical Study. The empirical studies are crucial for advancing our understanding and capabilities in detecting LLM-generated texts, as shown in Fig. 5 f). Recent work has examined the generalization and robustness of detection models, revealing notable challenges and insights. Tu et al.[290] observe a decline in RoBERTa-based detectors’ effectiveness over time through a month-long evaluation of ChatGPT’s long-form answers. Similarly, Pu et al.[291] show that detectors trained on medium-sized LLMs can generalize in zero-shot settings to larger models. Xu et al.[48] further find that detectors often overfit to dataset-specific artifacts due to insufficient data curation, yet are still capable of learning transferable features across domains and tasks. Sarvazyan et al.[292] report that in-domain fine-tuned detectors perform poorly when tested on outputs from different models, and MAGE[34] confirms this by evaluating 27 LLMs across 10 diverse datasets, showing that detectors frequently collapse under out-of-distribution settings. He et al.[294] present MGTBench, benchmarking 13 detectors on six LLMs and three datasets, and find that while model-based methods generalize better across datasets, metric-based ones transfer more effectively across models. They also highlight that even small amounts of training data suffice for competitive performance, though domain shifts still cause significant degradation. Other studies focus on paraphrase detection, highlighting threats to academic integrity.
Wahle et al.[30] demonstrate that humans struggle to detect paraphrased GPT-3 outputs, with only 53% accuracy, while also rating the paraphrases as highly fluent and coherent. Becker et al.[293] show that human-written paraphrases tend to be more diverse and difficult than machine-generated ones. Human evaluation further reveals consistent differences in writing style and content. Guo et al.[50] find that ChatGPT responses are more objective, formal, and detailed than human-written answers, while Chen et al.[295] discover that LLM-generated misinformation can be more difficult to detect than human-written misinformation, for both humans and LLM-based detectors like GPT-4 and LLaMA2.
In terms of attribution, Antoun et al.[296] conduct a large-scale study involving 50 LLMs and find an inverse correlation between classifier accuracy and model size—larger models are harder to attribute, especially when the classifier is trained on smaller models. Other studies propose representation- and feature-based attribution techniques. Munir et al.[313] identify subtle textual traces inherited from source models, while TopRoBERTa[314] incorporates topological data analysis to enhance RoBERTa’s ability to capture both semantic and structural patterns. Stylometric and psycholinguistic features are also leveraged: Uchendu et al.[168] use entropy and readability scores; GPT-who[307] adopts Uniform Information Density; and Wu et al.[308] apply proxy perplexity scores to attribute texts to models like LLaMA and OPT[337]. Soto et al.[309] propose a few-shot method that learns style features only from human-authored texts, avoiding reliance on generated samples. Finally, Sniffer[310] identifies model lineage by analyzing statistical discrepancies between model outputs, successfully linking derived models like Alpaca [338] to their predecessors such as ChatGPT and LLaMA.
3.1.3 Analysis
$\clubsuit$ Insights. Despite significant advances, most detection methods for LLM-generated text converge on a shared insight: LLMs produce outputs with detectable statistical regularities due to their decoding strategies and optimization objectives. These include over-regular syntax, repetitive phrasing, low grammatical variability, and coherent entity chains [325, 321, 329, 327]. Detection methods either exploit these patterns through white-box signals, such as token-level perplexity [321], log-probability curvature [323, 316], and cross-model comparisons [318, 317], or rely on black-box cues, including revision similarity and grammar correction [329, 328]. While white-box approaches offer higher precision, they require access to model internals or APIs [319], whereas black-box techniques are more flexible but often less robust. Several methods demonstrate strong zero-shot generalization by leveraging generation-intrinsic features rather than dataset-specific cues [322, 319]. However, trade-offs persist between interpretability and performance: interpretable methods like watermarking or edit-based signals are easy to explain but fragile, while statistical detectors offer higher accuracy at the cost of opacity [335, 326]. Notably, short-form content remains especially difficult to detect due to limited context, prompting recent interest in Positive-Unlabeled frameworks and style-based adaptation [324, 309]. Future detection systems may benefit from hybrid designs that combine white-box precision, black-box accessibility, and interpretability, while remaining resilient across languages, formats, and model variants.
$\clubsuit$ Recommendations. In light of the above analysis, several general strategies can be recommended for improving LAIM-generated text detection systems. To address the reliance on model internals, hybrid frameworks that combine white-box precision with black-box accessibility may offer a balance between accuracy and deployability. For instance, approximating internal signals (e.g., via top-K outputs or surrogate modeling) can expand the applicability of otherwise restricted methods. To enhance zero-shot generalization, it is beneficial to prioritize detection cues rooted in generation-intrinsic behaviors, such as token-level statistical regularities, rather than model-specific representations. For improved interpretability without sacrificing performance, a layered design that combines transparent surface-level heuristics with deeper statistical or neural backends can support both expert and non-expert use cases. Lastly, the persistent challenge of short form content calls for the integration of uncertainty-aware or semi-supervised learning techniques, such as Positive-Unlabeled modeling or few-shot adaptation, which can help maintain robustness under sparse or noisy inputs.
$\clubsuit$ Social Media. The widespread availability of LLMs has raised significant concerns about AI-generated content on social media platforms. Early works emphasized the challenges and initiated systematic efforts to detect and quantify this phenomenon. In late 2023, Lucas et al. [54] proposed the F3 framework, leveraging GPT-3.5’s generative capabilities not only to create synthetic disinformation but also to detect it using in-context, zero-shot reasoning techniques. Their findings highlighted that modern LLMs, if properly prompted, can outperform traditional fine-tuned detectors in identifying disinformation posts on platforms like Twitter and news outlets. Around the same time, Cui et al. [340] introduced the SAID benchmark, one of the first efforts to collect real-world AI-generated texts from social media platforms like Zhihu [341] and Quora [342]. SAID emphasized that detection in real-world settings is substantially harder than in simulated environments, and that user context information can greatly improve AI detection accuracy.
To address the multilingual and multi-platform gap, Macko et al.[66] introduced MultiSocial, a benchmark covering 22 languages and five platforms (e.g., Twitter, Telegram, Gab), showing that fine-tuned detectors outperform zero-shot methods on short, informal texts, though platform and language differences still pose challenges. Meanwhile, Sun et al.[67] conducted a large-scale study on Medium, Reddit, and Quora using the SM-D dataset, revealing a sharp rise in AI-generated content after late 2022, especially on Medium and Quora—and noting that such posts are often more objective, technical, and standardized than human-written ones.
Kumarage et al. [326] tackled the problem of detecting AI-generated tweets at a timeline level by developing a stylometric change point detection framework. Rather than detecting individual tweets, their method identifies shifts in writing style across a user’s posts, enabling forensic tracking of account takeovers by AI systems. Further complicating the landscape, Gameiro et al. [343] demonstrated that even state-of-the-art LLM detectors underperform in real-world settings, particularly when facing short news-like posts crafted by attackers who manipulate sampling parameters. Their findings stressed the urgent need for domain-specific benchmarks and cautioned against relying on existing detectors for operational deployment. Tuck and Verma expanded this line of inquiry by evaluating how censorship removal (e.g., uncensoring models like LLaMA3 and Qwen2) affects detectability [344]. They showed that uncensored models produce more human-like outputs, significantly undermining standard detection methods, and called for detectors capable of adapting to new domains and generation strategies. Finally, Ayoobi et al. [345] extended the concern to professional social networks like LinkedIn, where they introduced the SSTE (Section and Subsection Tag Embedding) method for detecting LLM-generated fake profiles. Their results demonstrated that even with limited training data, structured stylometric embeddings could robustly flag AI-crafted user profiles at registration time.
In conclusion, studies consistently show that detection accuracy drops significantly in real-world conditions [340, 343], where generation parameters are unknown and text is often shorter and noisier. Moreover, the rapid adoption of increasingly human-like uncensored models further weakens traditional detection strategies [344]. However, novel approaches, such as stylometric change point analysis [326], multilingual platform specific tuning [66], and in-context LLM-based detection [54], have shown promise in adapting to these challenges. Attribution and detection methods based on writing style rather than token probabilities offer improved generalization, while timeline and user level modeling introduce new avenues for forensic tracking.